OpenStudy (anonymous):

a) Given that cos x = sqrt3 - 1 find the value of cos 2x in the form a+b sqrt3 where a and b are integers b)Given that 2 cos(y+30) = sqrt 3 sin(y-30) find the value of k sqrt3

OpenStudy (anonymous):

Let ... √3.sin x + cos x = 1. Then, multiplying both sides by (1/2), (√3 /2) sin x + (1/2) cos x = 1/2 sin x · cos π/6 + cos x · sin π/6 = sin π/6 sin [ x + (π/6) ] = sin π/6 x + (π/6) = 2nπ ± π/6, ... n is an integer.

OpenStudy (anonymous):

that is square root

OpenStudy (anonymous):

Thank you the '.' represent multiplication right?